OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)?

OpenStudy (anonymous):

Find Slope. \[m = \frac{ y2 - y1 }{ x2 - x1 }\](-3, 2) and (-5, 8) \[m = \frac{ 8 - 2 }{ -5 -(-3) } = \frac{ 8-2 }{ -5+3 }=\frac{ 6 }{ -2 }=-3\]

OpenStudy (anonymous):

Now put into Point-Slope form. y - y1 = m(x - 1) m is slope Substitute x1 and y1 for the given point's coordinates. Slope of -3 (-3, 2) y - 2 = -3(x - (-3)) y - 2 = -3(x + 3) Now change into Slope-Intercept form.y = mx + b y - 2 = -3(x + 3) Distribute. y - 2 = -3x - 9 Add 2 to both sides. y = -3x - 7

OpenStudy (anonymous):

Sorry it took a while for me to finish up.

OpenStudy (anonymous):

Its ok can you help me with something else? @jazy

OpenStudy (anonymous):

Sure